### v1

# Towards gravimetry enhancement with squeezed states

Abstract
We investigate the estimation sensitivity of gravitational acceleration using squeezed probe states within a quantum metrology framework. In particular, we analyze how the squeezing phase, beyond its amplitudes, of the probes affects the attainable precision. We find that probes squeezed along the canonical phase-space quadrature can fail to achieve a quantum Fisher information (QFI) surpassing the shot-noise limit, regardless of the interaction time with the gravitational field. In contrast, position-momentum correlated input states with the squeezing amplitude can overcome this limit. Furthermore, we show that optimal sensitivity is attained through projective momentum measurements combined with a time-dependent adjustment of the squeezing phase. Our results are important to highlight the fundamental role of phase-engineered squeezing in experimental gravimetry protocols.
preprint: APS/123-QED
Introduction. The pursuit of novel protocols towards high-precision measurements of gravitational acceleration is crucial in the field of gravimetry [1, 2, 3, 4]. Experimentally, significant advances have been made in gravimeters employing several quantum techniques, including atomic interferometry [5, 6, 7], Bose–Einstein condensate (BEC) [4, 8, 9], ultracold neutrons [10], superconducting systems [11], and optomechanics systems  [12, 13]. Theoretically, improvements in the sensing of gravitational fields are pivotal in fundamental tests of general relativity theory [14, 15, 16, 17, 18] and also in the ongoing challenges to unify quantum mechanics and general relativity [19, 20, 21, 22]. Furthermore, from a technological perspective, ultrasensitive gravimeters are essential for developments in geological surveys and navigation, for instance [23, 24, 25].
Accurate estimation of physical quantities is crucial for theoretical developments and experimental breakthroughs. In estimation theory, the Cramér-Rao inequality, $\text{Var}(\hat{\alpha})\geq[n\mathcal{I}_{\alpha}(x)]^{-1/2}$, establishes an essential lower bound on the sensitivity (i.e., the standard deviation) with which an unknown parameter, $\alpha$, encoded in a conditional probability $p(x|\alpha)$ of observing the random value $x$, can be estimated [26, 27]. Here, $\hat{\alpha}$ denotes an unbiased estimator that maps the outcome $x$ to an estimate of the parameter $\alpha$, $n$ is the number of independent experiment trials, and $\mathcal{I}_{\alpha}(x)$ is the classical Fisher information (CFI). The CFI quantifies how changes in the parameter value $\alpha$ affect the probability distribution of the measured quantity $x$, through a particular choice of measurement strategy or a set of positive operator-valued measurements (POVM) [28, 29]. Given its dependence on the specific measurement strategy, the CFI can be optimized for all possible POVM [26, 28, 29], leading to an ultimate upper limit for the CFI, namely Quantum Fisher information (QFI), $\mathcal{F}_{\alpha}=\text{Tr}[\hat{\rho}_{\alpha}\hat{L}_{\alpha}]$, where $\hat{L}_{\alpha}$ is the symmetric logarithmic derivative defined as $\partial_{\alpha}\hat{\rho}_{\alpha}=(\hat{L}_{\alpha}\hat{\rho}_{\alpha}+\hat{\rho}_{\alpha}\hat{L}_{\alpha})/2$ [30, 28, 29]. The quantum parameter estimation theory establishes the QFI as a fundamental precision limit for estimating the parameter $\alpha$, bounding the classical Fisher information, i.e., $\mathcal{F}_{\alpha}\geq\mathcal{I}_{\alpha}$ and giving rise to the quantum Cramér–Rao inequality, which defines the minimal achievable estimation variance employing a quantum state [31, 28, 29, 30, 32].
Quantum metrology aims to explore the features of quantum probes such as entanglement [33, 28, 29], coherence [34], non-classicality [35, 36], contextuality [37], and squeezing [38, 39, 40, 41], to overcome the precision of sensing unknown parameters compared to their classical counterparts [32, 28, 29], leading to a series of experimental advances in this field [42, 43]. In particular, since their original conception [44, 45] and pioneering proposal for use in laser interferometric gravitational wave detectors [38], squeezed states have found applications in a wide range of gravimetry protocols [46]. Despite the recognized advantage of employing squeezed probe states in gravimetry protocols, most applications have mainly focused on the role of the squeezing amplitude or have considered situations where the squeezing phase is aligned with a canonical phase-space quadrature, such as position or momentum.
In this Letter, we investigate the boost in the estimation of the gravitational acceleration, $g$, by considering input squeezed states endowed with a squeezing phase that differs from the canonical quadratures of position and momentum. Specifically, we analyze how the gravitational acceleration is dynamically encoded in both the amplitude and phase of the squeezed probe and how different phases of squeeze affect the QFI. We demonstrate that by choosing an appropriate phase of the squeezed input state, one can create a non-null correlation between the position and momentum quadratures, thereby significantly increasing the QFI and improving the sensitivity in estimating the gravitational acceleration, which overcomes vacuum probe states, regardless of the free-falling time under the gravitational field. Additionally, we show that a projective momentum measurement can saturate the QFI bound and therefore achieve this optimal sensitivity. Notably, standard quadrature squeezing along position or momentum does not always yield optimal results, highlighting the importance of the fine-tuning of the squeezing phase in quantum metrology protocols.
The model. We consider the dynamics of a quantum particle with inertial mass, $m$, free falling in a non-relativistic gravitational potential, $U(z)=mgz$, where $z$ is the distance from the floor and $g$ denotes the gravitational acceleration. Experimentally, this scenario can be feasible by employing ultra-cold atoms that freely fall from a vertical tower in the gravitational potential [9, 10]. The system is initialized in a Gaussian state,
$\rho_{0}(z_{0},z_{0}^{\prime})=\mathcal{N}_{0}\exp\left[-\frac{(z_{0}^{2}+z_{0}^{\prime 2})}{4\sigma^{2}}+\frac{i\gamma(z_{0}^{2}-z_{0}^{\prime 2})}{4\sigma^{2}}\right],$ (1)
that formally corresponds to a squeezed vacuum state, where $r$ and $\theta$ represent the squeezing amplitude and phase, respectively. $\mathcal{N}_{0}=1/(\sqrt{2\pi}\sigma)$ denotes the normalization constant, $\sigma=\sigma_{0}[\cosh(2r)-\sinh(2r)\cos(2\theta)]^{1/2}$ is the position uncertainty at initial time, $\sigma_{0}$ the standard deviation of the vacuum state, and $\gamma=\sinh(2r)\sin(2\theta)$ quantifies the position-momentum correlation [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The phase $\theta$ sets the quadrature along which squeezing is applied and determines the orientation of the uncertainty ellipse in phase space (See Supplemental Material (SM) [51]). For instance, when $\theta=0$ the squeezing is along the position quadrature, $\sigma=\sigma_{0}/e^{2r}$ (reduction in position uncertainty), and for $\theta=\pi/2$, the squeezing is along the momentum quadrature, $\sigma=\sigma_{0}e^{2r}$ (increasing in position uncertainty). For intermediate values of $\theta$, the squeezed input state (1) exhibits a non-null correlation between the position and momentum quadratures [57]. Recent studies have highlighted the role of such position-momentum correlations as a valuable resource in metrological tasks [58, 59, 60, 61]. Position-momentum correlations are a key resource for universal continuous-variable quantum computation (CVQC) [62], as established by a recent resource theory [63].
Fig. 1: QFI for estimation of $g$ as a function of the evolution time $\tau/\tau_{0}$ for different amplitudes and phases of the squeezed input state. Panels (a-c) correspond to the short-time evolution regime, while panels (d-f) depict the dynamics at the long-time limit. Curves represent squeezing intensities $r=0.4$ (solid black line), $r=0.5$ (dashed red line), $r=0.6$ (dotted blue line), for distinct squeezing phases $\theta$ of the squeezed input state (1) and vacuum state $r=0$ (dot-dashed gray line). Both panels illustrate that the quantum advantage in estimating gravitational acceleration with squeezed probe states initialized in canonical directions, such as position ($\theta=0$) and momentum ($\theta=\pi/2$), is limited to a certain time [shaded area in Fig. 1(a, e)]. In contrast, probes with position-momentum correlation offer a consistent improvement in the QFI over any vacuum state, regardless of the evolution time [shaded area in Fig. 1(c, f)].
Gravitational acceleration estimation. Using the path integral formalism, the evolution of the input squeezed state (1) under a uniform gravitational field can be analytically determined (See SM for more details [51]). Additionally, the QFI concerning the parameter $g$ is given by (details in SM [51])
$\begin{split}\mathcal{F}_{g}(\tau,r,\theta)&=\frac{\sinh^{2}(2r)\sin^{2}(2\theta)+1}{4\sigma_{0}^{2}[\cosh(2r)-\sinh(2r)\cos(2\theta)]}\tau^{4}\\ &\quad+\frac{2\tau_{0}\sinh(2r)\sin(2\theta)}{\sigma_{0}^{2}}\tau^{3}\\ &\quad+\frac{4\tau_{0}^{2}[\cosh(2r)-\sinh(2r)\cos(2\theta)]}{\sigma_{0}^{2}}\tau^{2},\end{split}$ (2)
where $\tau_{0}=m\sigma_{0}^{2}/\hbar$ is a constant with time dimension. In the particular case where $r=0$ (i.e., $\gamma=0$), the QFI (2) becomes $\mathcal{F}_{g}^{\text{vac}}(\tau)=\tau^{4}/4\sigma^{2}_{0}+4\tau_{0}^{2}\tau^{2}/\sigma^{2}_{0}$, consistent with the result obtained in Ref. [64] for an input vacuum state. Some conclusions can be drawn from Eq. (2). First, the QFI depends on both amplitude and phase of the squeezing. Consequently, it allows for the possibility of setting up different squeezing phases of the input probe, rather than restricting squeezing to the conventional quadratures by fixing $\theta=0$ or $\pi/2$, to enhance the sensitivity of gravitational acceleration compared to the input vacuum state. Moreover, since the energy of the input probe state is independent of $\theta$, possible advantages in the QFI by changing the squeezing phase come without any additional energy cost. Second, the QFI is insensitive to the intensity of the gravitational acceleration, meaning that the capacity to estimate the $g$ does not change with the strength of this parameter. Third, for short ($\tau\ll\tau_{0}$) and long ($\tau\gg\tau_{0}$) interaction times, the QFI (2) evolves into $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\theta)\approx[\sigma^{2}/\sigma_{0}^{2}]\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ and $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\theta)\approx[\sigma_{0}^{2}/\sigma^{2}+\gamma^{2}]\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$, respectively. This characteristic enables fine-tuning the squeezing intensity $r$ and phase $\theta$ of the input probe to enhance beyond that of the vacuum state, thereby increasing the sensitivity to $g$ over specific interaction times.
Figure 1 presents the QFI (2) for estimating the gravitational acceleration parameter $g$ as a function of the dimensionless interaction time $\tau/\tau_{0}$, considering different input probe states. In all plots, we set the following parameters $m=1$, $\sigma_{0}=1$, $\hbar=1$ to focus only on the general trends and features of the dynamics. Panels (a-c) correspond to the dynamic of $\mathcal{F}_{g}(\tau,r,\theta)$ for short interaction times, whereas panels (d-f) illustrate the behavior for long times. The curves represent the squeezing intensities $r=0.4$ (solid black line), $r=0.5$ (dashed red line), $r=0.6$ (dotted blue line), for distinct squeezing phases of the input state (1), and the vacuum state $r=0$ (dot-dashed gray line). For the input states squeezed in momentum ($\theta=\pi/2$), the QFI surpasses that of the vacuum probe only in the short-time regime, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\pi/2)\approx e^{2r}\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [shaded area in Fig. 1(a)], the input vacuum state being more sensitive in estimating the gravitational acceleration $g$ for long times, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\pi/2)\approx e^{-2r}\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(d)]. Conversely, for probes squeezed in position ($\theta=0$), the QFI enhancement over the input vacuum state occurs exclusively in the long time limit, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,0)\approx e^{2r}\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [shaded area in Fig. 1(e)], with the input vacuum state more sensitive for short times, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,0)\approx e^{-2r}\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(b)]. Interestingly, probe states initially endowed with position-momentum correlations ($\theta=\pi/4$) achieve a QFI greater than that of the vacuum state in both the short, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\pi/4)\approx\cosh(2r)\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(c)], and long, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\pi/4)\approx(\sinh^{2}2r+\operatorname{sech}2r)\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(f)], interaction-time regimes, respectively. This result highlights the fact that the quantum advantage in estimating gravitational acceleration using squeezed probes initialized in the canonical position ($\theta=0$) and momentum ($\theta=\pi/2$) direction is limited to a certain time domain. In contrast, squeezed probes endowed with initial position-momentum correlation can achieve a consistent improvement in the QFI over the vacuum state, regardless of the interaction time.
Relative quantum Fisher information. To quantify the possible enhancement from squeezed states in estimating the gravitational acceleration beyond the vacuum probe, we define the relative QFI (RQFI)
$\mathcal{Q}(\tau,r,\theta)=\frac{\mathcal{F}_{g}(\tau,r,\theta)}{\mathcal{F}_{g}^{\text{vac}}(\tau)}.$ (3)
For $\mathcal{Q}>1$, the squeezed probe provides a precision advantage over the vacuum state for estimating gravitational acceleration. The case $\mathcal{Q}<1$ represents no advantage of employing squeezed states, while $\mathcal{Q}=1$ corresponds to the equivalence between both the vacuum and the squeezed probe in the estimation of $g$.
Figure 2(a) illustrates the time evolution of $\mathcal{Q}(\tau,r,\theta)$ for three distinct squeezing phases, where the behaviors in short ($\tau\ll\tau_{0}$) and asymptotic regimes ($\tau\gg\tau_{0}$) are summarized in Table 1. We consider probe states with squeezing applied along the position quadrature (solid black line, $\theta=0$), the momentum quadrature (dashed red line, $\theta=\pi/2$), and an intermediate value corresponding to the position-momentum correlated state (dotted blue line, $\theta=\pi/4$). These results highlight the fact that the sensitivity of the gravitational acceleration strongly depends on the phase of the input squeezed probe. For a squeezed probe in position ($\theta=0$), the RQFI monotonically increases with time, reaching its asymptotic limit $e^{2r}$, whereas for a probe squeezed in momentum $\theta=\pi/2$ it decays toward $e^{-2r}$, indicating a loss of metrological advantage $Q<1$. As a result, the squeezing advantage $\mathcal{Q}>1$ (shaded area) provided by probe states with squeezing in position or momentum crucially depends on the interaction time. Conversely, by employing the position-momentum correlated state ($\theta=\pi/4$), with the same quantity of squeezing, it always exhibits a squeezing advantage to estimate the gravitational acceleration $g$ in relation to the vacuum probe, regardless of the evolution time.
Table 1: $\mathcal{Q}(\tau,r,\theta)$ for different squeezing phases in the short and asymptotic time regimes.
$\theta$ short-time ($\tau\ll\tau_{0}$) asymptotic ($\tau\gg\tau_{0}$)
0 $e^{-2r}$ $e^{2r}$
$\pi/2$ $e^{2r}$ $e^{-2r}$
$\pi/4$ $\cosh(2r)$ $\cosh^{2}(2r)$
Despite our choice of $\theta=\pi/4$ in Fig. 2(a), there are other position-momentum correlated state with a QFI surpassing that of the vacuum state. Figure 2(b) displays a contour plot of $\mathcal{Q}$ as a function of both time $\tau$ and the squeezing phase $\theta$, allowing us to explore the continuous variation of metrological performance in different squeezing orientations. In particular, the results reveal that the enhanced sensitivity for all times ($\mathcal{Q}>1$) is not always limited to the specific case $\theta=\pi/4$, but can be achieved across a wide range of angles, highlights that optimal metrological gain for estimation of gravitational acceleration does not depend solely on the squeezing strength but critically on the phase-space orientation of the squeezed quadratures.
Fig. 2: The Relative Quantum Fisher Information (RQFI) $\mathcal{Q}$, with values $\mathcal{Q}>1$ implies an advantage using squeezed states, $\mathcal{Q}<1$ indicates no gain, and $\mathcal{Q}=1$ denotes equal performance with vacuum states in the estimation of $g$. Panel (a) illustrates the temporal evolution of $\mathcal{Q}$ for different squeezing directions in phase space. The solid black line corresponds to squeezing applied along the canonical position quadrature direction ($\theta=0$), the dashed red line represents squeezing along the momentum quadrature ($\theta=\pi/2$), and the dotted blue line depicts a probe exhibiting position-momentum correlations, with the squeezing direction chosen at an intermediate phase of $\theta=\pi/4$. In (b), we display a contour plot of $\mathcal{Q}$ as a function of time and squeezing phase $\theta$. It highlights that multiple directions beyond $\theta=\pi/4$ yield $\mathcal{Q}>1$, demonstrating the broad flexibility in optimizing squeezing orientation for enhanced metrological performance. In all plots, we fixed the squeezing amplitude at $r=0.5$.
Measurement schemes. Although the QFI establishes the ultimate precision bound for an unbiased estimator, $\hat{\alpha}$, the optimal measurement required to achieve this bound is often difficult or impractical [28, 65, 66]. This is because it requires a projective measurement associated with POVM elements given by projectors onto the eigenstates of the logarithmic derivative operator $\hat{L}_{\alpha}$. To identify measurement schemes capable of approaching the bound imposed by the QFI, we focus on strategies that can be readily described within the Gaussian formalism [67]. We consider local Gaussian projective measurements, which can be efficiently modeled by updating the covariance matrix as $\widetilde{\boldsymbol{\sigma}}=\boldsymbol{\sigma}+\boldsymbol{\sigma}_{m}$ [67, 66], where the matrix $\boldsymbol{\sigma}_{m}=\text{Diag}[s,s^{-1}]/2$ characterizes the measurement and is parametrized by $s$. This framework encompasses projective position measurements ($s\rightarrow 0$), projective momentum measurements ($s\rightarrow\infty$), as well as homodyne and heterodyne detection ($s=1$). The exact CFI, $\mathcal{I}_{g}(\tau,r,\theta,s)$, that quantifies the information extracted by a given measurement strategy [67, 66] can be analytically determined in terms of the parameter $s$ (see supplementary material [51]). By scanning the values of $s$, it is possible to identify which measurement strategy is most effective in achieving the ultimate bound, where the CFI approaches the QFI.
Fig. 3: Behavior of the ratio $\mathcal{R}(\tau,r,\theta,s)=\mathcal{I}_{g}(\tau,r,\theta,s)/\mathcal{F}_{g}(\tau,r,\theta)$ between the CFI and the QFI bound for different measurement schemes, as a function of the squeezing phase $\theta$ and the evolution time $\tau$. Here, the squeezing parameter $r=0.5$ is kept fixed across all panels and the measurement setting $s$ varies from one panel to another. Panel (a) shows the case $s\rightarrow\infty$ (projective momentum measurement), (b) corresponds to $s\rightarrow 0$ (projective position measurement), and (c) illustrates the heterodyne scheme with $s=1$. Note that only the momentum measurement strategy can saturate the QFI bound, yielding the maximum possible information $\mathcal{R}=1$, yellow region in Fig. 3(a)]. Yet, the optimal choice of squeezing phase is time-dependent, and squeezing along canonical quadratures, position ($\theta=0$) or momentum ($\theta=\pi/2$), does not always provide the best sensitivity at all times.
In Fig. 3, we show the ratio $\mathcal{R}(\tau,r,\theta,s)=\mathcal{I}_{g}(\tau,r,\theta,s)/\mathcal{F}_{g}(\tau,r,\theta)$ between the CFI, $\mathcal{I}_{g}(\tau,r,\theta,s)$, and the QFI bound, $\mathcal{F}_{g}(\tau,r,\theta)$, for different measurement strategies. The ratio $\mathcal{R}$ quantifies how closely a given measurement approaches the ultimate precision bound set by the QFI. Here, the squeezing parameter $r=0.5$ is kept fixed in all panels, and the measurement setting $s$ varies in each. We analyze how different squeezing phases $\theta$ of the squeezed input state can be adjusted over time to reach the QFI limit. The limiting situation $s\rightarrow\infty$, which denotes an ideal momentum measurement, is shown in Fig. 3(a). This approach can theoretically reach the highest sensitivity allowed by quantum physics in this domain, since the CFI can saturate the QFI bound, i.e., $\mathcal{I}_{g}=\mathcal{F}_{g}$ ($\mathcal{R}=1$) [lighter yellow region, in Fig. 3(a)]. However, this optimality depends on the appropriate choice of the squeezing phase, which changes with time $\tau$. The complementary scenario $s\rightarrow 0$, corresponding to a projective position measurement, is displayed in Fig. 3(b). Unlike the momentum example, the projective position measurement does not saturate the QFI constraint, $\mathcal{R}\approx 0.4$ [yellow region in Fig. 3(b)]. The heterodyne detection scheme with $s=1$ is shown in Figure 3(c). This approach also fails to reach the QFI limit $\mathcal{R}\approx 0.5$, yellow region in Fig. 3(c)], highlighting the importance of searching for appropriate measurement strategies in quantum metrology. Interestingly, optimal sensitivity is not always achieved by canonical quadrature squeezing, i.e., squeezed probe with phase along position ($\theta=0$) or momentum ($\theta=\pi/2$). In addition, the time dependence of the optimal phase $\theta_{\text{opt}}(\tau)$ suggests that adaptive or dynamically controlled squeezed strategies may be needed to approach the quantum limit over different time scales consistently.
Conclusion. In this work, we have employed squeezed probe states to estimate the gravitational acceleration parameter. We analyze how both the amplitude and phase of the squeezing influence the sensitivity of the estimation. Our results show that the metrological advantage in gravitational acceleration estimation, compared to a vacuum probe state, is not solely determined by the squeezing strength but crucially depends on the squeezing phase, i.e, it depends on the direction of the applied squeezing in the probe state. Specifically, while probes squeezed along the canonical quadratures may fail to surpass the shot-noise limit, position–momentum correlated states with the same squeezing magnitude can exhibit a quantum Fisher information exceeding that of the vacuum state. Moreover, we identified optimal measurement strategies and squeezing phases that allow the classical Fisher information to reach the quantum limit. These findings deepen our understanding of how the phase of squeezed probes can be exploited to enhance precision in quantum gravimetry protocols.
Acknowledgments. O.R.A. acknowledges the Graduate Program in Physics at Federal University of Piauí. L.S.M. acknowledges support from the National Institute of Science and Technology on the National Institute of Photonics (INFO) CNPq - INCT grant 409174/2024-6. J. F. G. S. acknowledges CNPq Grant No. 420549/2023-4, Fundect Grant No. 83/026.973/2024, and Universidade Federal da Grande Dourados for support. C.H.S.V. acknowledges the São Paulo Research Foundation (FAPESP) Grant No. 2023/13362-0 and Grant No. 2025/14546-2 for financial support and the Southern University of Science and Technology (SUSTech) for providing the workspace.